Janet​, an experienced shipping​ clerk, can fill a certain order in 5 hours. Felipe​, a new​ clerk, needs 8 hours to do the same job. Working​ together, how long will it take them to fill the​ order?
Answer by fractalier(6550)

We set this up this way
t/5 + t/8 = 1
Now multiply by 40 and get
8t + 5t = 40
13t = 40
t = 40/13 hours